### v2

# Hidden Breakthroughs in Language Model Training

Abstract
Loss curves are smooth during most of model training, so visible discontinuities stand out as possible conceptual breakthroughs. Studying these breakthroughs enables a deeper understanding of learning dynamics, but only when they are properly identified. This paper argues that similar breakthroughs occur frequently throughout training but they are obscured by a loss metric that collapses all variation into a single scalar. To find these hidden transitions, we introduce POLCA, a method for decomposing changes in loss along arbitrary bases of the low-rank training subspace. We use our method to identify clusters of samples that share similar changes in loss during training, disaggregating the overall loss into that of smaller groups of conceptually similar data. We validate our method on synthetic arithmetic and natural language tasks, showing that POLCA recovers clusters that represent interpretable breakthroughs in the model’s capabilities. We demonstrate the promise of these hidden phase transitions as a tool for unsupervised interpretability.
1 Introduction
As large language models train, various internal structures develop during abrupt phase transitions. These sudden drops in loss reveal the formation of mechanisms for in-context learning [25], natural language grammar [6], hierarchical generalization [21], and many other concepts [19, 16, 27, 1]. However, the loss curve as a whole remains stubbornly smooth. As a result, these momentary conceptual breakthroughs are treated as isolated curiosities, while the majority of training behavior is considered predictable.
This work shows that in fact, a model undergoes many breakthroughs during training, most which are concealed when averaging all data into a single loss curve. Instead, we divide up the loss curve in two different ways to find hidden breakthroughs. First, we disaggregate the aggregate loss into losses on individual examples. By clustering the individual loss curves, we identify subsets of data that experience synchronized changes in loss, implying that they rely on the same conceptual breakthrough. However, any individual example might benefit from multiple breakthroughs; such an example may undergo changes synchronized with different data subsets at different times. Furthermore, distinct concepts might appear simultaneously, erroneously merging their data clusters. This means we may have to identify multiple separate breakthroughs which affect the loss curve of a single example.
To disentangle these effects for a single sample, we separate the optimization space into specific gradient directions. When the loss changes during training, it is the result of movement across all parameters in a high-dimensional space. We decompose this loss change from an exact trajectory in the full-rank parameter space into a collection of movements along each dimension. By analyzing these loss curves along specific basis vectors, we identify conceptual breakthroughs that rely on particular directions of movement. The latter analysis permits further granularity in clustering data, as final performance on an individual example may rely on multiple conceptual breakthroughs, each corresponding to a particular linear direction in training. In summary:
We show that some learned concepts can be identified by clustering exact loss, while others cannot (Section 4.2).
Using POLCA, we extend our cluster analysis to identify conceptual breakthroughs in a restricted gradient subspace that are obscured in the exact loss curves. We automatically identify specific concepts learned during breakthroughs in both synthetic (Section 4) and natural language settings (Section 5).
2 Background: How much can we learn from learning dynamics?
Various loss phase transitions have been interpreted as conceptual breakthroughs. But why expect additional interpretable breakthroughs to underlie periods of undifferentiated, gradual model improvement? Our approach is justified by the nature of the loss surface’s complexity, illlustrated by Figure 1, in which a smooth curve emerges by eliding phase transitions from each dimension.
Why expect multiple phase transitions?
A very early phase transition is to be expected early in training after a brief memorization stage [30]. In this sense, the most celebrated breakthroughs are not the first phase transitions in their training curves. Some breakthroughs even depend on earlier breakthroughs, as observed in synthetic tasks [1] and in grammar acquisition [6]. If one concept depends on another, each must appear at a different timestep, requiring multiple phase transitions.
Multiple phase transitions can also come from differences in gradient scale along different directions. Ma et al. [18] even attributed the early edge-of-stability phase transition [13, 8] to multiscale structure of the loss surface and, furthermore, noted that this multiscale structure emerges at the range where models become singular: their loss lacks a quadratic approximation in terms of model parameters, creating conditions for phase transitions under Singular Learning Theory [33, 34, 32]. They argued that this structure is the product of both nonuniform data and nonconvex objectives, respectively justifying the disaggregation and decomposition which we apply to interpret training dynamics.
Why disaggregate the aggregate loss?
We track learning on training datapoints and subpopulations, rather than the whole training set, because relevant skills can be acquired at different rates [3, 7]. Individual samples thus exhibit changes in loss out of line with the monotonic average trend [36, 29]. In full-batch gradient descent, Cohen et al. [8] identified non-monotonicity arising from oscillation about the maximum Hessian eigenvector. Rosenfeld and Risteski [29] demonstrated that these oscillations occur across different axes for different samples and identified the primary cause: surprisingly human-interpretable semantic features. Even when the loss seems stable, performance can oscillate on edge cases until the model develops relevant capabilities [28, 4]. We hypothesize that oscillation represents competing skills that are relevant to different subsets of data. To test this hypothesis, and to interpret the meaning of these directions, we disaggregate the loss into clusters with similar dynamics.
Why decompose the exact loss?
Our POLCA decomposition addresses what Michaud et al. [20] called polygenic scaling effects—samples which combine multiple skills and therefore exhibit breakthroughs at multiple scales. If we assume that a specific skill is enabled by movement along a particular skill’s basis vector, then the loss change attributed to movement along that vector will accelerate at the moment the skill is acquired, for every sample that requires that skill. In this manner, the sample transitions from early to late dynamics through a basis-specific loss phase transition. In other words, by monitoring changes in directions corresponding to specific skills, we support the speculation of Nanda et al. [23] that “phase transitions are everywhere.”
Why is linear decomposition sufficient?
In practice, a conceptual breakthrough might not occur in a single direction that persists throughout training. However, there is an abundance of evidence that linear bases of the low-rank [12] training subspace are conceptually meaningful. In the late stages of training, loss is convex on the line connecting a pair of checkpoints [10] if those checkpoints express similar capabilities [14] and mechanisms [17]. If a pair of high-dimensional models lack this linear connection, they still connect nonlinearly [9]; however, while parameter settings sampled from their linear connections improve broadly on the capabilities of the original models, those sampled from their nonlinear connections are less robust than the originals [14, ref Appendix D]. These observations suggest that linear decomposition should preserve meaningful conceptual features on the loss surface, and our experiments show that the resulting directions are interpretable in practice.
3 Methods
The key to our approach is the separate consideration of how each example’s datapoint loss changes throughout training. We contrast this individualized metric with aggregated loss across an entire dataset. Using the datapoint loss, we can cluster individual examples on the basis of their loss , change in loss , or magnitude of change during training.
Our objective is to decompose the loss itself into specific directions in the weight space, motivated by several considerations: First, while we have moved from an aggregated loss metric to a more granular datapoint loss metric, we are still only considering breakthroughs that are general enough to be perceived in loss curves. Second, an individual datapoint may benefit from a variety of conceptual breakthroughs, but will not be clustered on the breakthroughs individually. Finally, once we have identified a subset of the data as benefiting from a particular conceptual breakthrough, decomposing into individual weight directions allows us to locate where in the weights the breakthrough occurs and to thereby identify the mechanism involved.
Next we break this loss down by directional movement during training, allowing us to discover breakthroughs that are specific to a given direction. Our procedure follows three steps: (1) select a basis, (2) decompose the loss along that basis to highlight particular learning events, (3) cluster datapoints according to their shared learning events.
3.1 Finding the basis
To decompose the loss, we first require an interpretable orthogonal basis. We efficiently compute the eigenvectors of the Hessian matrix with CoLA [26] and use them to construct a restricted training subspace. We posit this basis to be interpretable because each basis vector expresses a high gradient covariance and therefore represents a potential decision boundary. While we select this basis for interpretability, our approach can use any arbitrary choice of basis.
This basis is constructed as shown in Algorithm 1. Given intermediate training checkpoints and a number of eigenvectors to compute at each checkpoint, we seek a low rank -dimensional subspace which captures most of the movement during optimization [12]. We construct this basis iteratively, starting with : at each checkpoint , we take checkpoint weights and project their loss Hessian onto the nullspace of . From the resulting projection, we append the top eigenvectors to . We compute the eigenvectors using Hessian-vector products [11] to avoid explicitly constructing the full Hessian matrix. The resulting basis is designed to include directions of highest curvature at each checkpoint so that it will capture synchronized loss behavior throughout training. Note that the very top eigenvectors are likely to reflect local oscillation, rather than conceptually meaningful long-term movement [31], but as we continue to add to the low rank basis, we include more directions of long-term stable movement. We discard the oscillatory directions which do not provide an overall decrease in loss over the course of training according to POLCA.
3.2 Decomposing the loss with Projection Oriented Loss Change Allocation (POLCA)
To decompose the loss along our basis, we propose a modified version of Loss Change Allocation (LCA) [15]. LCA is a tool for analyzing changes in aggregated loss on dataset between two checkpoints. The output of LCA is the empirical loss change between a pair of checkpoints which can be attributed to the motion of each individual weight unit. Given two consecutive checkpoints with parameters and , LCA reformulates the change in loss as its first-order Taylor approximation, a sum of the loss changes attributed to the movement of each individual model parameter :
The POLCA decomposition differs from LCA in three key ways. First, we do not restrict each direction to correspond to a single unit , instead permitting an arbitrary orthonormal basis vector to replace the axis-aligned basis vectors in LCA; we project onto this basis vector using the dot product . Second, we are interested in changes in the loss on each individual example , not the entire dataset . These first two modifications provide the following reformulation of LCA.
The third key difference is that we must use a second-order approximation because this basis is constructed explicitly from the Hessian eigenvectors. To understand why this choice of basis requires a second-order approximation, recall that each basis vector is an eigenvector of the Hessian matrix at some timestep , where is chosen because it has the largest eigenvalue over the whole dataset. If we assume that the top eigenvectors of the aggregate Hessian maintain high curvature at other points in training and on individual datapoints, then the scaling factor in the second-order Taylor term will be very large even at the datapoint level. Limiting the approximation to only the first order term gives poor guarantees on error, as the second-order term could be expected to dominate. We find that empirically, the difference between the first and second-order values is small (Appendix F), but compute the second-order approximation to guarantee a better estimate.
Exact computation of the second-order term would be intractable, requiring computation of the top eigenvalues/vectors for each individual datapoint . Instead, we can approximate it by substituting the true eigenvalue, denoted , with the curvature of the individual loss in the direction , i.e. . If the aggregate Hessian eigenvector is close to the span of the top eigenvectors of the datapoint-specific Hessian for , this provides a reasonable estimate while reducing calculation to a single Hessian-vector product per eigenvector. We therefore approximate the basis projection of the datapoint Hessian as derived in Appendix A.
Equipped with this second-order approximation of the datapoint Hessian’s projection onto our basis, we account for the high curvature and possible domination by the higher order term by modifying Equation 4 into the second-order Taylor expansion using the approximation from Equation 7.
3.3 Clustering the loss
POLCA, above, provides curves that show how a decomposed loss changes with respect to each training example. We assume that if several examples show similarly timed loss changes in the same direction, they likely rely on the same conceptual breakthroughs or learning events; therefore, they are likely to share a required skill. We cluster POLCA training histories to recover these skill groups. For each datapoint , we compute the total cumulative change in loss along each basis vector by summing over the previous POLCA values. We denote this sum the projected loss .
We obtain 1d projected loss trajectories for breakthrough clustering by computing at every time . As our clustering algorithm, we choose Hierarchical Density-Based Spatial Clustering of Applications with Noise (HDBSCAN) [5] because it distinguishes cluster outliers and discovers clusters with variable density (i.e., similarly shaped curves that lay far apart in their metrics).
4 Arithmetic language modeling
We validate our POLCA clustering method in a synthetic setting using an arithmetic addition task. Our clusters reflect categorical concepts within the data, even when those concepts are not discoverable by clustering directly on loss curves. Specifically, if we cluster on exact loss curves we recover digit positions, but if we cluster on POLCA curves we also recover the skill of “carrying” a digit.
4.1 Experiments
Data
Our synthetic experiments use data from the arithmetic addition setting in Chen et al. [7], where the model is trained to compute the sum of two 3-digit numbers. As shown in Appendix Figure 7 and Chen et al. [7], the skills corresponding to the digits have different loss curves, so we will easily recover the digit skill categories by clustering exact loss curves. Unlike our source material, we also consider an additional skill: arithmetic carries to the output token (Figure 3). Digit-specific addition skills lead to clearly distinguishable loss curves, whereas carrying skills are not clear from the exact loss (Appendix Figure 8), but become clear on the decomposed gradient basis.
Setup details We train a 3-layer (9m parameter) Transformer model with embedding dimension 512, 4 attention heads per layer, and an MLP dimension of 2048 [22]. We choose this model size to align with prior work Olsson et al. [24] and to maximize the granularity at which we can feasibly compute the POLCA values. For a validation set with 1250 data points and 5000 output tokens, we compute the loss and POLCA values for each token at intervals of 20 iterations throughout training. We choose the training steps between each POLCA computation to achieve as fine-grained analysis as possible without exploding the compute time. We compute the POLCA basis using the eigenvectors of the Hessian estimated using a 1250 data point sample of the training set as detailed in Algorithm 1. We compute one new basis vector every 200 iterations for a total of 50 basis vectors. We provide further ablations on the decomposition strategy and choice of POLCA basis in Appendix D and E.
Clustering
To ensure that we are investigating directions where the model is learning on average, we only consider the basis vectors for which the mean projected loss decreases. Then for each remaining basis vector, we remove all of the tokens for which the decomposed loss increases. We use HDBSCAN to cluster the remaining tokens, discarding the tokens it marks as outliers. Through this process, we find subpopulations of the data that have similar projected loss trajectories.
4.2 Results
Comparison to the exact loss
In our clustering experiments on arithmetic addition skills, we first consider whether directional decomposition is necessary for identifying individual concepts. As a baseline, we therefore cluster tokens solely on their exact loss curves for successive timesteps, rather than estimating the loss decomposed along a low rank basis.
According to the HDBSCAN loss clustering results in Figure 3, we can recover—to a substantial degree—the digit skill by clustering only on the exact loss, likely because the digits have very different loss trajectories. However, as shown in Figure 3 and Appendix Table 3, we cannot recover clusters that are homogenous with respect to the carrying from the exact loss alone.
Recovering concepts with POLCA clustering
Because exact loss clustering failed to recover the carrying skill, we instead cluster on each basis vector’s projected loss using the POLCA decomposition. The projected loss value (Equation 10) represents the cumulative loss change of attributed to movement along basis vector . By clustering the projected loss trajectories, we find that the top 3 basis vectors produce homogeneous clusters corresponding to both digit and carrying skills for each of the digits (Figure 4 and Appendix Figure 9), so POLCA clustering is able to recover subtler skills, like carrying, that are challenging to reconstruct from the exact loss curves alone.
Figure 4 shows the POLCA breakthrough clustering for the first three basis vectors. Along these directions, certain data subpopulations have changes in the projected loss (Figure 4(a) and 4(d)) that do not occur as visibly in their full loss curves (Figure 4(b), 4(e)). We conclude that arithmetic carries rely on breakthroughs along specific dimensions during training, but these breakthroughs may be elided in the exact loss curve.
5 English language modeling
We apply our approach to a real-world causal language modeling task and show that POLCA breakthrough clustering recovers interpretable conceptual skills in the natural language setting.
5.1 Experiments
For the natural language modeling setting, we use the English Wikipedia dataset [35] from March 2022. We use the same POLCA setup as in the arithmetic addition setting. We train a 3-layer (40m parameter) transformer model with embedding dimension 512, 4 attention heads, and an MLP dimension of 2048 [22]. We compute the loss and POLCA values for each token on a validation set of 12600 output tokens. We analyze intermediate checkpoints at intervals of 200 iterations throughout training. We apply POLCA to the basis derived from the eigenvectors of the Hessian estimated using a 1000 data point sample of the training set as detailed in Algorithm 1 with . We compute a new basis vector every 750 iterations.
Similarly to the arithmetic setting, we only analyze the POLCA trajectories along the directions for which the decomposed loss decreases on average. We also discard any trajectories for which the decomposed per-token loss increases. After clustering the remaining token trajectories with HDBSCAN, we discard any marked as outliers before analysis.
| Cluster | Label | Contexts closest to medoid |
[…]
| 3 | Comma after parenthetical phrase | in most countries. The largest episcopal conferences are those of India and China, followed by the Philippines, in 1888 by the Swedish physicist Johannes Rydberg, theoretically by Niels Bohr in 1913, klas Athletic F.C. is a football club based in Leagrave, in Luton, |
| Cluster | Label | Contexts closest to medoid |
[…]
5.2 Results
On language models, POLCA clustering again reveals hidden breakthroughs along each basis vector. To analyze the learned skills represented by each cluster, we look for syntactic and lexical patterns shared by the cluster data. After tagging tokens by part-of-speech (POS) using spacy, we label each cluster by computing its most common token and POS tag sequences. We confirm these automatically assigned cluster labels by examining the top 10 contexts closest to the medoid of the cluster.
We report a selection of clusters from specific basis vectors in Figure 6 (see Appendix H for more examples). Using POLCA clustering on projected loss trajectories for our basis vectors, we find token subpopulations that correspond to various grammatical constructions and have breakthroughs in the loss projected onto that basis. For instance, we show a cluster corresponding to predicting < to> and < from> after the first clause in a sentence in Figure 6a(6ii). We also observe clusters whose projected loss trajectories move in opposing directions along certain basis vectors. For instance, in Figure 6b(6i) and 6b(6ii), cluster 1 contains appositive noun phrases and cluster 2 has noun phrases with similar syntactic patterns that are not appositive (such as list items). These clusters visibly mirror each other’s movement in direction, if not in magnitude—decreases in loss are generally larger than the opposing cluster’s increases in loss which mirror them.
Figure 6 shows that despite their smooth exact loss curves, POLCA clusters have sudden changes in their decomposed loss curves at different points during training. Clusters from the exact loss curves, by contrast, do not reveal breakthroughs except very early in training (Appendix Figure 10). We conclude that POLCA reveals breakthroughs in the decomposed loss that are obscured in the exact loss. Through clustering, POLCA can explain how different skills are learned during language model training.
6 Conclusions
This work introduces POLCA clustering, a method to identify learned skills from decomposed loss trajectories. POLCA decomposes the loss on two levels: individual datapoints and specific directions in the weight space. We use this decomposition to discover clusters that share breakthroughs obscured by loss metrics. In language modeling and synthetic settings, these clusters recover interpretable skills which appear to emerge at particular moments during training.
These are promising findings for meaningfully interpreting large models. By recovering breakthroughs in identifiable skills, we support the hypothesis that high-dimensional learning typically entails a series of phase transitions at various scales. When a phase transition appears in training, it suggests a naturally discrete category; the model either knows the concept or doesn’t know it, with little middle ground. Humans think in categorical concepts, so they are far more interpretable than the continuous data interpolations that appear in much of learning theory.
Limitations
Our method of constructing a basis is inspired by the existing literature on training in restricted subspaces, but represents an obvious site of improvement. The top eigenvectors of the Hessian, like the axis-aligned basis, could represent many concepts in superposition. Therefore, some non-orthogonal basis might represent interpretable concepts more cleanly than our orthogonal basis, though it would no longer provide a low-rank decomposition. Furthermore, our basis is constructed by local curvature and then filtered to favor directions of long-term movement; other bases may favor long-term movement by construction. In general, we consider the ideal basis to be an open question.
We use small models so that we can have very fine-grained checkpoints for POLCA computation. Scaling this work to larger models may require using a basis that is less computationally expensive to compute than Hessian eigenvectors, but our results from Appendix E indicate that this is likely possible with limited impact on the cluster quality.
Acknowledgements
This work was informed by helpful conversations with Nikhil Vyas, Nicholas Lourie, Mike Lepori, and Ekdeep Singh Lubana. This material is based upon work supported by the National Science Foundation Graduate Research Fellowship under Grant No. DGE 2140743. Any opinion, findings, and conclusions or recommendations expressed in this material are those of the authors(s) and do not necessarily reflect the views of the National Science Foundation. This work was enabled in part by a gift from the Chan Zuckerberg Initiative Foundation to establish the Kempner Institute for the Study of Natural and Artificial Intelligence.
Appendix A Derivation of approximate second-order term
We can approximate the difference between the gradient at time and as
If we assume to also be an eigenvector of the datapoint Hessians , we can apply a similar argument for the gradient on the datapoint level.
Note that the assumption above (of matching Hessian eigenvectors between data points and their aggregate) is unlikely to be correct. If this assumption is violated, then the scaling factor in the following second-order Taylor term will be minuscule on the datapoint level. In practice, we have found that the second-order term has limited impact at the datapoint level (see Appendix F), but we nonetheless use it to improve our approximation. Then we may approximate it as:
Appendix B Hyperparameters
In the tables below, we provide the hyperparameters used during training of the models in the synthetic arithmetic and language modeling settings. We use NVIDIA H100 80GB HBM3 GPUs for our experiments and run each training run on a single GPU.
| HYPERPARAMETER | VALUE |
| Number of Parameters | 9475594 |
| Iterations | 10000 |
| Epochs | 1 |
| Batch size | 64 |
| Number of training tokens | 2560000 |
| Optimizer | AdamW |
| Learning rate | 1e-5 |
| Weight decay | 0.1 |
| Betas | |
| LR Schedule |
| HYPERPARAMETER | VALUE |
| Number of Parameters | 40274737 |
| Iterations | 14000 |
| Epochs | 1 |
| Batch size | 64 |
| Number of training tokens | 114688000 |
| Optimizer | AdamW |
| Learning rate | 1e-5 |
| Weight decay | 0.1 |
| Betas | |
| LR Schedule |
Appendix C Exact loss trajectories for the digit and carry skills
Appendix D Decomposition strategy comparison
We investigate whether POLCA is required for decomposing the loss. We test three strategies:
Loss: Exact loss
Fisher information: We approximate the empirical Fisher Information as as in [2]. For each basis vector , the Fisher Information projected onto is .
POLCA: We compute the projected loss (Equation 9).
While using the projected gradient is less computationally intensive than computing the entire POLCA term, the results from Table 3 demonstrate that clustering on solely the Fisher information does not yield sufficiently homogeneous carry clusters.
| Decomposition strategy | Top vector skill homogeneities | ||
| 1 | 2 | 3 | |
| Loss | 0.514 | - | |
| Fisher information | 0.664 | 0.643 | 0.637 |
| POLCA | 0.973 | 0.946 | 0.903 |
Appendix E POLCA basis comparison
We test ablated bases to analyze the effect of basis choice on the POLCA breakthrough clustering. To do so, we compute the maximum carry skill homogeneities over all of the clusters when performing POLCA breakthrough clustering. We use the following bases:
Random orthonormal: randomly sampled orthonormal vectors
Random shuffled Hessian: basis computed using Algorithm 1, but randomly shuffling the model checkpoints
Top Hessian eigenvectors: basis computed using Algorithm 1
We find in Table 4 that these ablations result in only slightly less homogeneous carry clusters, indicating that different bases can be used for larger experiments to trade off between compute and interpretability.
| POLCA basis | Top vector skill homogeneities | ||
| 1 | 2 | 3 | |
| Random orthonormal | 0.902 | 0.858 | 0.838 |
| Random shuffled Hessian | 0.856 | 0.852 | 0.834 |
| Top Hessian eigenvectors | 0.973 | 0.946 | 0.903 |
Appendix F Second versus first order POLCA approximation
| Cosine similarity | L2 norm |
| 5.4891 E-4 | 0.99987 |
Appendix G Additional arithmetic language modeling clusters
Appendix H Additional natural language clusters
| Cluster | Label | Contexts closest to medoid |
[…]
| Cluster | Label | Contexts closest to medoid |
[…]
| Cluster | Label | Contexts closest to medoid |
[…]

### v4

@@
 Loss curves are smooth during most of model training, so visible discontinuities stand out as possible conceptual breakthroughs.
-Studying these breakthroughs enables a deeper understanding of learning dynamics, but only when they are properly identified.
-This paper argues that similar breakthroughs occur frequently throughout training but they are obscured by a loss metric that collapses all variation into a single scalar.
+These breakthroughs enable a deeper understanding of the model’s concept structure, but only when they are properly identified.
+This paper argues that similar breakthroughs occur frequently throughout training, but they are obscured by a loss metric that collapses all variation into a single scalar.
 To find these hidden transitions, we introduce POLCA, a method for decomposing changes in loss along arbitrary bases of the low-rank training subspace.
 We use our method to identify clusters of samples that share similar changes in loss during training, disaggregating the overall loss into that of smaller groups of conceptually similar data.
-We validate our method on synthetic arithmetic and natural language tasks, showing that POLCA recovers clusters that represent interpretable breakthroughs in the model’s capabilities.
-We demonstrate the promise of these hidden phase transitions as a tool for unsupervised interpretability.
+We validate our method on synthetic arithmetic and English language modeling, showing that POLCA recovers clusters that represent interpretable breakthroughs in the model’s capabilities.
+We demonstrate the promise of these hidden breakthroughs as a tool for unsupervised interpretability.
 1 Introduction
-As large language models train, various internal structures develop during abrupt phase transitions.
-These sudden drops in loss reveal the formation of mechanisms for in-context learning [25], natural language grammar [6], hierarchical generalization [21], and many other concepts [19, 16, 27, 1].
+As large language models train, various internal structures develop during abrupt breakthroughs.
+These sudden drops in loss reveal the formation of mechanisms for in-context learning (Olsson et al., 2022b), natural language grammar (Chen et al., 2024a), hierarchical generalization (Murty et al., 2023), and many other concepts (McGrath et al., 2022; Lovering et al., 2022; Power et al., 2022; Abbe et al., 2021).
 However, the loss curve as a whole remains stubbornly smooth.
 As a result, these momentary conceptual breakthroughs are treated as isolated curiosities, while the majority of training behavior is considered predictable.
-This work shows that in fact, a model undergoes many breakthroughs during training, most which are concealed when averaging all data into a single loss curve.
-Instead, we divide up the loss curve in two different ways to find hidden breakthroughs.
+These breakthroughs—often colloquially termed phase transitions (Olsson et al., 2022a; Chen et al., 2024a; Murty et al., 2023)—are extremely consequential for our understanding of neural networks.
+Phase transitions represent critical periods of learning, so they offer key insights for training and optimization.
+For instance, introducing noisy data or changing the optimizer during a phase transition can significantly reduce the downstream performance of a model (Achille et al., 2017; Chen et al., 2024a).
+Their timing is used as evidence the role of specific mechanisms in model capabilities (Zhong et al., 2023; Olsson et al., 2022b; Chen et al., 2024a).
+If we could find more of these momentary training events, it could expand our understanding even further.
+Prior work identifies breakthroughs through a top-down approach by measuring the training dynamics of a predefined concept or skill and searching for sudden changes.
+We instead propose a bottom-up unsupervised method for finding breakthroughs by grouping data points that have similar training behavior.
+This data-centric approach can be used to inform optimization choices such as data selection or learning rate scheduling.
+Like other bottom-up interpretability methods such as SAEs, PCA, and transcoders, our method seeks concepts that are used naturally by the model, rather than imposing an assumed structure onto learning and representation.
+This work shows that in fact, a model undergoes many breakthroughs during training, but most are concealed when averaging all data into a single loss curve.
+Instead of averaging, we divide up the loss curve in two different ways to find hidden breakthroughs.
 First, we disaggregate the aggregate loss into losses on individual examples.
@@
 Furthermore, distinct concepts might appear simultaneously, erroneously merging their data clusters.
-This means we may have to identify multiple separate breakthroughs which affect the loss curve of a single example.
+To recover the underlying learned concepts, we may have to identify multiple separate breakthroughs which affect the loss curve of a single example.
 To disentangle these effects for a single sample, we separate the optimization space into specific gradient directions.
@@
 We show that some learned concepts can be identified by clustering exact loss, while others cannot (Section 4.2).
-Using POLCA, we extend our cluster analysis to identify conceptual breakthroughs in a restricted gradient subspace that are obscured in the exact loss curves.
+Using POLCA, we extend our cluster analysis to identify these hidden conceptual breakthroughs obscured in the exact loss curves.
 We automatically identify specific concepts learned during breakthroughs in both synthetic (Section 4) and natural language settings (Section 5).
 2 Background: How much can we learn from learning dynamics?
-Various loss phase transitions have been interpreted as conceptual breakthroughs.
+Various loss breakthroughs have been interpreted as learning specific concepts.
 But why expect additional interpretable breakthroughs to underlie periods of undifferentiated, gradual model improvement?
-Our approach is justified by the nature of the loss surface’s complexity, illlustrated by Figure 1, in which a smooth curve emerges by eliding phase transitions from each dimension.
-Why expect multiple phase transitions?
-A very early phase transition is to be expected early in training after a brief memorization stage [30].
-In this sense, the most celebrated breakthroughs are not the first phase transitions in their training curves.
-Some breakthroughs even depend on earlier breakthroughs, as observed in synthetic tasks [1] and in grammar acquisition [6].
-If one concept depends on another, each must appear at a different timestep, requiring multiple phase transitions.
-Multiple phase transitions can also come from differences in gradient scale along different directions.
+Our approach is justified by the nature of the loss surface’s complexity, illustrated by Figure 1, in which a smooth curve emerges by eliding breakthroughs from each dimension.
+Why expect multiple breakthroughs?
+A very early phase transition is to be expected early in training after a brief memorization stage (Shwartz-Ziv & Tishby, 2017).
+In this sense, the most celebrated breakthroughs—those signaling the formation of induction heads or arithmetic grokking—are not the first drops in their training curves.
+Some breakthroughs even depend on earlier breakthroughs, as observed in synthetic tasks (Abbe et al., 2021) and in grammar acquisition (Chen et al., 2024a).
+If one concept depends on another, each must appear at a different timestep, requiring multiple breakthroughs.
+Furthermore, as shown by Saxe et al.
+(2019), summing many phase transitions can result in a smooth curve, supporting our hypothesis that these breakthroughs can appear in stable regions of the loss curve.
+Multiple breakthroughs can also come from differences in gradient scale along different directions.
 Ma et al.
-[18] even attributed the early edge-of-stability phase transition [13, 8] to multiscale structure of the loss surface and, furthermore, noted that this multiscale structure emerges at the range where models become singular: their loss lacks a quadratic approximation in terms of model parameters, creating conditions for phase transitions under Singular Learning Theory [33, 34, 32].
+(2022) even attributed the early edge-of-stability phase transition (Jastrzȩbski et al., 2020; Cohen et al., 2022) to multiscale structure of the loss surface and, furthermore, noted that this multiscale structure emerges at the range where models become singular: their loss lacks a quadratic approximation in terms of model parameters, creating conditions for breakthroughs under Singular Learning Theory (Watanabe, 2010; Wei et al., 2020; Wang et al., 2024).
 They argued that this structure is the product of both nonuniform data and nonconvex objectives, respectively justifying the disaggregation and decomposition which we apply to interpret training dynamics.
 Why disaggregate the aggregate loss?
-We track learning on training datapoints and subpopulations, rather than the whole training set, because relevant skills can be acquired at different rates [3, 7].
-Individual samples thus exhibit changes in loss out of line with the monotonic average trend [36, 29].
+We track learning on training datapoints and subpopulations, rather than the whole training set, because relevant skills can be acquired at different rates (Arora & Goyal, 2023; Chen et al., 2024b).
+Individual samples thus exhibit changes in loss out of line with the monotonic average trend (Xia et al., 2023; Rosenfeld & Risteski, 2024).
 In full-batch gradient descent, Cohen et al.
-[8] identified non-monotonicity arising from oscillation about the maximum Hessian eigenvector.
-Rosenfeld and Risteski [29] demonstrated that these oscillations occur across different axes for different samples and identified the primary cause: surprisingly human-interpretable semantic features.
-Even when the loss seems stable, performance can oscillate on edge cases until the model develops relevant capabilities [28, 4].
+(2022) identified non-monotonicity arising from oscillation about the maximum Hessian eigenvector.
+Rosenfeld & Risteski (2024) demonstrated that these oscillations occur across different axes for different samples and identified the primary cause: surprisingly human-interpretable semantic features.
+Even when the loss seems stable, performance can oscillate on edge cases until the model develops relevant capabilities (Qin et al., 2024; Bhaskar et al., 2024).
 We hypothesize that oscillation represents competing skills that are relevant to different subsets of data.
@@
 Why decompose the exact loss?
-Our POLCA decomposition addresses what Michaud et al.
-[20] called polygenic scaling effects—samples which combine multiple skills and therefore exhibit breakthroughs at multiple scales.
-If we assume that a specific skill is enabled by movement along a particular skill’s basis vector, then the loss change attributed to movement along that vector will accelerate at the moment the skill is acquired, for every sample that requires that skill.
-In this manner, the sample transitions from early to late dynamics through a basis-specific loss phase transition.
+Michaud et al.
+(2024) analyzed the scaling behavior of models with respect to individual tokens and identified a limitation of token-wise analysis of breakthroughs, which they called polygenic scaling effects—samples which combine multiple skills and therefore exhibit breakthroughs at multiple scales.
+Our POLCA decomposition directly addresses this limitation by decomposing the loss of each token along multiple basis vectors.
+If we assume that a specific skill is enabled by movement along that skill’s basis vector, then the loss change attributed to movement along that vector will accelerate at the moment the skill is acquired, for every sample that requires that skill.
+In this manner, the sample transitions from early to late dynamics through a basis-specific loss breakthrough.
 In other words, by monitoring changes in directions corresponding to specific skills, we support the speculation of Nanda et al.
-[23] that “phase transitions are everywhere.”
+(2023) that “phase transitions are everywhere.”
 Why is linear decomposition sufficient?
 In practice, a conceptual breakthrough might not occur in a single direction that persists throughout training.
-However, there is an abundance of evidence that linear bases of the low-rank [12] training subspace are conceptually meaningful.
-In the late stages of training, loss is convex on the line connecting a pair of checkpoints [10] if those checkpoints express similar capabilities [14] and mechanisms [17].
-If a pair of high-dimensional models lack this linear connection, they still connect nonlinearly [9]; however, while parameter settings sampled from their linear connections improve broadly on the capabilities of the original models, those sampled from their nonlinear connections are less robust than the originals [14, ref Appendix D].
+However, there is abundant evidence that linear bases of the low-rank (Gur-Ari et al., 2018) training subspace are conceptually meaningful.
+In the late stages of training, loss is convex on the line connecting a pair of checkpoints (Frankle et al., 2020) if those checkpoints express similar capabilities (Juneja et al., 2023) and mechanisms (Lubana et al., 2023).
+If a pair of high-dimensional models lack this linear connection, they still connect nonlinearly (Draxler et al., 2019); however, while parameter settings sampled from their linear connections improve broadly on the capabilities of the original models, those sampled from their nonlinear connections are less robust than the originals (Juneja et al., 2023, ref Appendix G).
 These observations suggest that linear decomposition should preserve meaningful conceptual features on the loss surface, and our experiments show that the resulting directions are interpretable in practice.
 3 Methods
-The key to our approach is the separate consideration of how each example’s datapoint loss changes throughout training.
+The key to our approach is the separate consideration of each example’s datapoint loss changes throughout training.
 We contrast this individualized metric with aggregated loss across an entire dataset.
-Using the datapoint loss, we can cluster individual examples on the basis of their loss , change in loss , or magnitude of change during training.
-Our objective is to decompose the loss itself into specific directions in the weight space, motivated by several considerations: First, while we have moved from an aggregated loss metric to a more granular datapoint loss metric, we are still only considering breakthroughs that are general enough to be perceived in loss curves.
+Using the datapoint loss, we can cluster individual example on the basis of its loss , change in loss , or magnitude of change during training.
+We next decompose the loss itself into specific directions in the weight space, motivated by several considerations: First, while we have moved from an aggregated loss metric to a more granular datapoint loss metric, we are still only considering breakthroughs that are general enough to be perceived in loss curves.
 Second, an individual datapoint may benefit from a variety of conceptual breakthroughs, but will not be clustered on the breakthroughs individually.
 Finally, once we have identified a subset of the data as benefiting from a particular conceptual breakthrough, decomposing into individual weight directions allows us to locate where in the weights the breakthrough occurs and to thereby identify the mechanism involved.
-Next we break this loss down by directional movement during training, allowing us to discover breakthroughs that are specific to a given direction.
+When we break the exact loss curve down by directional movement during training, the resulting decomposed loss will reveal breakthroughs that are specific to a given direction.
 Our procedure follows three steps: (1) select a basis, (2) decompose the loss along that basis to highlight particular learning events, (3) cluster datapoints according to their shared learning events.
@@
 To decompose the loss, we first require an interpretable orthogonal basis.
-We efficiently compute the eigenvectors of the Hessian matrix with CoLA [26] and use them to construct a restricted training subspace.
-We posit this basis to be interpretable because each basis vector expresses a high gradient covariance and therefore represents a potential decision boundary.
-While we select this basis for interpretability, our approach can use any arbitrary choice of basis.
-This basis is constructed as shown in Algorithm 1.
-Given intermediate training checkpoints and a number of eigenvectors to compute at each checkpoint, we seek a low rank -dimensional subspace which captures most of the movement during optimization [12].
+We efficiently compute the eigenvectors of the Hessian matrix using CoLA (Potapczynski et al., 2023) to construct a restricted training subspace.
+We expect this basis to be interpretable because each basis vector captures a large gradient covariance and therefore represents a potential decision boundary.
+We select this basis for interpretability, but our approach can use an arbitrary choice of basis—for example, one which targets a particular use case or efficiently leverages optimizer preconditioner values.
+The basis is constructed as shown in Algorithm 1.
+Given intermediate training checkpoints and a number of eigenvectors to compute at each checkpoint, we seek a low rank -dimensional subspace which captures most of the movement during optimization (Gur-Ari et al., 2018).
 We construct this basis iteratively, starting with : at each checkpoint , we take checkpoint weights and project their loss Hessian onto the nullspace of .
 From the resulting projection, we append the top eigenvectors to .
-We compute the eigenvectors using Hessian-vector products [11] to avoid explicitly constructing the full Hessian matrix.
+We compute the eigenvectors using Hessian-vector products Golmant et al.
+(2018) to avoid explicitly constructing the full Hessian matrix.
 The resulting basis is designed to include directions of highest curvature at each checkpoint so that it will capture synchronized loss behavior throughout training.
-Note that the very top eigenvectors are likely to reflect local oscillation, rather than conceptually meaningful long-term movement [31], but as we continue to add to the low rank basis, we include more directions of long-term stable movement.
-We discard the oscillatory directions which do not provide an overall decrease in loss over the course of training according to POLCA.
-3.2 Decomposing the loss with Projection Oriented Loss Change Allocation (POLCA)
-To decompose the loss along our basis, we propose a modified version of Loss Change Allocation (LCA) [15].
+Note that the very top eigenvectors are likely to reflect local oscillation, rather than conceptually meaningful long-term movement (Song et al., 2024), but as we continue to add to the low rank basis, we include more directions of long-term stable movement.
+We discard the oscillatory directions which do not provide an overall decrease in loss over the course of training according to POLCA by removing the directions with an increase in the mean projected loss from checkpoint to .
+In this manner, we first construct a basis based on local information, then filter out directions that do not represent long-term movement.
+This construction finds local high-curvature directions that may be important for breakthroughs in the intermediate stages of training while ensuring that the basis does not overfit to local oscillations.
+3.2 Decomposing the loss with POLCA
+To decompose the loss along our basis, we propose a modified version of Loss Change Allocation (LCA) (Lan et al., 2020) which we call Projection-Oriented Loss Change Allocation (POLCA).
 LCA is a tool for analyzing changes in aggregated loss on dataset between two checkpoints.
@@
 Second, we are interested in changes in the loss on each individual example , not the entire dataset .
-These first two modifications provide the following reformulation of LCA.
-The third key difference is that we must use a second-order approximation because this basis is constructed explicitly from the Hessian eigenvectors.
-To understand why this choice of basis requires a second-order approximation, recall that each basis vector is an eigenvector of the Hessian matrix at some timestep , where is chosen because it has the largest eigenvalue over the whole dataset.
+These first two modifications provide the first-order POLCA decomposition:
+The third key difference is that we use a second-order approximation because this basis is constructed explicitly from the Hessian eigenvectors.
+To understand why this choice of basis warrants a second-order approximation, recall that each basis vector is an eigenvector of the Hessian matrix at some timestep , where is chosen because it has the largest eigenvalue over the whole dataset.
 If we assume that the top eigenvectors of the aggregate Hessian maintain high curvature at other points in training and on individual datapoints, then the scaling factor in the second-order Taylor term will be very large even at the datapoint level.
 Limiting the approximation to only the first order term gives poor guarantees on error, as the second-order term could be expected to dominate.
-We find that empirically, the difference between the first and second-order values is small (Appendix F), but compute the second-order approximation to guarantee a better estimate.
+Although empirically the difference between the first and second-order values is small (see Appendix I), we nonetheless guarantee a better estimate due to lower Lagrange error bounds by computing the second-order approximation below.
 Exact computation of the second-order term would be intractable, requiring computation of the top eigenvalues/vectors for each individual datapoint .
@@
 If the aggregate Hessian eigenvector is close to the span of the top eigenvectors of the datapoint-specific Hessian for , this provides a reasonable estimate while reducing calculation to a single Hessian-vector product per eigenvector.
-We therefore approximate the basis projection of the datapoint Hessian as derived in Appendix A.
+We therefore approximate the basis projection of the datapoint Hessian as derived in Appendix C.
 Equipped with this second-order approximation of the datapoint Hessian’s projection onto our basis, we account for the high curvature and possible domination by the higher order term by modifying Equation 2 into the second-order Taylor expansion using the approximation from Equation 5.
+We can compute this second-order term with limited additional computational complexity by keeping track of the eigenvalues for each Hessian eigenvector and the aggregate gradient at each checkpoint.
 3.3 Clustering the loss
 POLCA, above, provides curves that show how a decomposed loss changes with respect to each training example.
-We assume that if several examples show similarly timed loss changes in the same direction, they likely rely on the same conceptual breakthroughs or learning events; therefore, they are likely to share a required skill.
+We assume that if several examples show similarly timed loss changes in the same direction, they likely rely on the same conceptual breakthroughs or learning events; therefore, they are likely to share a required skill, or specific capability needed for a given task.
 We cluster POLCA training histories to recover these skill groups.
@@
 We obtain 1d projected loss trajectories for breakthrough clustering by computing at every time .
-As our clustering algorithm, we choose Hierarchical Density-Based Spatial Clustering of Applications with Noise (HDBSCAN) [5] because it distinguishes cluster outliers and discovers clusters with variable density (i.e., similarly shaped curves that lay far apart in their metrics).
+We cluster trajectories using Hierarchical Density-Based Spatial Clustering of Applications with Noise (HDBSCAN) Campello et al.
+(2013) because it distinguishes cluster outliers and discovers clusters with variable density (i.e., similarly shaped curves that lay far apart in their metrics).
+We cluster the trajectories for each basis vector separately to ensure that the clustering can capture multiple skills per token.
+3.4 Defining and identifying hidden breakthroughs
+We use POLCA to recover hidden breakthroughs in training, so we must quantify whether clustered trajectories are distinguished by breakthroughs.
+We use the formulation defined by Chen et al.
+(2024b) to compute the start of a breakthrough in a given function for a given datapoint and basis vector :
+Here, approximates the maximum point of acceleration of in .
+can be either the projected loss for a given basis vector or the exact loss .
+We define a hidden breakthrough as a breakthrough that occurs in the flat region of the exact loss curve.
+That is, if we set a threshold beyond which the exact loss curve is flat, then a given set has a hidden breakthrough in a metric if the expected value of the start of breakthroughs in is greater than :
 4 Arithmetic language modeling
@@
 Our clusters reflect categorical concepts within the data, even when those concepts are not discoverable by clustering directly on loss curves.
-Specifically, if we cluster on exact loss curves we recover digit positions, but if we cluster on POLCA curves we also recover the skill of “carrying” a digit.
+Specifically, if we cluster on exact loss curves, we recover digit positions—but if we instead cluster on POLCA curves, we also recover the skill of “carrying” a digit.
 4.1 Experiments
@@
 Our synthetic experiments use data from the arithmetic addition setting in Chen et al.
-[7], where the model is trained to compute the sum of two 3-digit numbers.
+(2024b), where the model is trained to compute the sum of two 3-digit numbers.
+This setting has 4 skills corresponding to each of the digits in the output sum.
+Note that the digit in the 1000s place is always a <0> or <1> token since the two input summants are 3 digits long.
 As shown in Appendix Figure 7 and Chen et al.
-[7], the skills corresponding to the digits have different loss curves, so we will easily recover the digit skill categories by clustering exact loss curves.
+(2024b), the skills corresponding to the digits have different loss curves, so we will easily recover the digit skill categories by clustering exact loss curves.
 Unlike our source material, we also consider an additional skill: arithmetic carries to the output token (Figure 3).
-Digit-specific addition skills lead to clearly distinguishable loss curves, whereas carrying skills are not clear from the exact loss (Appendix Figure 8), but become clear on the decomposed gradient basis.
-Setup details We train a 3-layer (9m parameter) Transformer model with embedding dimension 512, 4 attention heads per layer, and an MLP dimension of 2048 [22].
-We choose this model size to align with prior work Olsson et al.
-[24] and to maximize the granularity at which we can feasibly compute the POLCA values.
-For a validation set with 1250 data points and 5000 output tokens, we compute the loss and POLCA values for each token at intervals of 20 iterations throughout training.
-We choose the training steps between each POLCA computation to achieve as fine-grained analysis as possible without exploding the compute time.
-We compute the POLCA basis using the eigenvectors of the Hessian estimated using a 1250 data point sample of the training set as detailed in Algorithm 1.
-We compute one new basis vector every 200 iterations for a total of 50 basis vectors.
-We provide further ablations on the decomposition strategy and choice of POLCA basis in Appendix D and E.
+This skill corresponds to the case where instead of simply adding the two tokens at the corresponding digit in the input, the model must also carry a from the previous digit.
+Digit-specific addition skills lead to clearly distinguishable exact loss curves, whereas carrying skills do not (Appendix Figure 8)—but carries will become clear on our decomposed gradient basis.
+We provide additional skill and labeling details in Appendix E.1.
+Setup details We train a 3-layer (9 million parameter) Transformer model with embedding dimension 512, 4 attention heads per layer, and an MLP dimension of 2048, following prior work (Olsson et al., 2022a).
+We study a validation set with 1250 data points and 5000 output tokens throughout training.
+We compute the loss and POLCA values for each token at each interval of 20 training steps.
+The POLCA basis uses the eigenvectors of the Hessian, estimated using a 1250 data point sample of the training set as detailed in Algorithm 1.
+We compute one new basis vector every 200 training steps, for a total of 50 basis vectors.
+We provide further ablations on the decomposition strategy and choice of POLCA basis in Appendices G and H.
+We train the model for steps, but trim the x-axis of the plots at to better display the breakthroughs.
 Clustering
+As described in Section 3.1, some of the top Hessian directions may represent directions of oscillation (and not learning) during training.
 To ensure that we are investigating directions where the model is learning on average, we only consider the basis vectors for which the mean projected loss decreases.
-Then for each remaining basis vector, we remove all of the tokens for which the decomposed loss increases.
+Then for each remaining basis vector, we remove all of the tokens for which the decomposed loss increases, therefore retaining only tokens which rely on the vector during training.
+This removes 2360.8 out of 5000 tokens on average—suggesting that any given direction is irrelevant to learning many examples.
 We use HDBSCAN to cluster the remaining tokens, discarding the tokens it marks as outliers.
@@
 4.2 Results
+| Exact loss | 0.514 | 0.0 | |||||||
+| Change in exact loss | 0.524 | 0.0 | |||||||
+| LCA (Lan et al., 2020) | 0.792 | 0.019 | |||||||
+| POLCA | 0.973 | 0.355 |
 Comparison to the exact loss
-In our clustering experiments on arithmetic addition skills, we first consider whether directional decomposition is necessary for identifying individual concepts.
-As a baseline, we therefore cluster tokens solely on their exact loss curves for successive timesteps, rather than estimating the loss decomposed along a low rank basis.
-According to the HDBSCAN loss clustering results in Figure 3, we can recover—to a substantial degree—the digit skill by clustering only on the exact loss, likely because the digits have very different loss trajectories.
-However, as shown in Figure 3 and Appendix Table 3, we cannot recover clusters that are homogenous with respect to the carrying from the exact loss alone.
+In our clustering experiments on arithmetic addition skills, we first consider whether decomposition is necessary for identifying conceptual skills.
+As a baseline, we therefore cluster tokens solely on their exact loss curves, rather than their decomposed loss curves.
+According to the exact loss clustering results in Figure 3, we can recover—to a substantial degree—the digit skill by clustering only on the exact loss, likely because the digits have very different loss trajectories.
+However, as shown in Figure 3 and Table 1, we cannot recover clusters that are homogenous with respect to the carrying skill from the exact loss alone.
 Recovering concepts with POLCA clustering
-Because exact loss clustering failed to recover the carrying skill, we instead cluster on each basis vector’s projected loss using the POLCA decomposition.
+Because exact loss clustering failed to recover the carrying skill, we will recover it with a different clustering method.
+Instead of exact loss, we will cluster on each basis vector’s projected loss using the POLCA decomposition.
 The projected loss value (Equation 8) represents the cumulative loss change of attributed to movement along basis vector .
-By clustering the projected loss trajectories, we find that the top 3 basis vectors produce homogeneous clusters corresponding to both digit and carrying skills for each of the digits (Figure 4 and Appendix Figure 9), so POLCA clustering is able to recover subtler skills, like carrying, that are challenging to reconstruct from the exact loss curves alone.
-Figure 4 shows the POLCA breakthrough clustering for the first three basis vectors.
-Along these directions, certain data subpopulations have changes in the projected loss (Figure 4(a) and 4(d)) that do not occur as visibly in their full loss curves (Figure 4(b), 4(e)).
+By clustering the projected loss trajectories, we find that the top 2 basis vectors produce two easily-described and near-homogeneous clusters: one which contains examples of the 1000s place digit and one which contains examples of arithmetic carrying for all digits (Figure 4 Appendix Figure 9).
+We therefore determine that POLCA clustering recovers subtler skills—like carrying—that are challenging to reconstruct from the exact loss or parameter-aligned LCA curves alone (Table 1).
+In Appendix Table 9, we also use Equation 10 to compare the fraction of clusters with hidden breakthroughs past step (with ), where the mean exact loss plateaus (Appendix Figure 7), and find in Table 1 that POLCA discovers the highest fraction of clusters with hidden breakthroughs.
+Figure 4 shows POLCA clustering for the first two basis vectors.
+Along these basis vectors, certain data subpopulations show changes in the projected loss (Figures 4(a) and 4(d)) that do not occur as visibly in their smoother exact loss curves (Figures 4(b) and 4(e)).
 We conclude that arithmetic carries rely on breakthroughs along specific dimensions during training, but these breakthroughs may be elided in the exact loss curve.
@@
 We apply our approach to a real-world causal language modeling task and show that POLCA breakthrough clustering recovers interpretable conceptual skills in the natural language setting.
-5.1 Experiments
-For the natural language modeling setting, we use the English Wikipedia dataset [35] from March 2022.
-We use the same POLCA setup as in the arithmetic addition setting.
-We train a 3-layer (40m parameter) transformer model with embedding dimension 512, 4 attention heads, and an MLP dimension of 2048 [22].
-We compute the loss and POLCA values for each token on a validation set of 12600 output tokens.
-We analyze intermediate checkpoints at intervals of 200 iterations throughout training.
-We apply POLCA to the basis derived from the eigenvectors of the Hessian estimated using a 1000 data point sample of the training set as detailed in Algorithm 1 with .
-We compute a new basis vector every 750 iterations.
-Similarly to the arithmetic setting, we only analyze the POLCA trajectories along the directions for which the decomposed loss decreases on average.
-We also discard any trajectories for which the decomposed per-token loss increases.
-After clustering the remaining token trajectories with HDBSCAN, we discard any marked as outliers before analysis.
 | Cluster | Label | Contexts closest to medoid |
@@
 […]
+5.1 Experiments
+For the natural language modeling setting, we use the English Wikipedia dataset (Wikimedia Foundation, 2022) from March 2022 to train a 3-layer (40m parameter) model.
+We use the same POLCA setup as in the arithmetic addition setting (see Appendix E.2 for details).
+As in the arithmetic setting, we only consider directions where the projected aggregate loss overall decreases to filter out directions of oscillation.
+We also discard the datapoint-specific POLCA trajectories along the directions where the token’s decomposed loss increases, which removes an average 6655.5 out of 12600 tokens per direction.
+After clustering the remaining token trajectories, we discard HDBSCAN-marked outliers.
+Automatic labeling.
+To analyze the concepts represented by each cluster, we look for syntactic and lexical patterns shared by the cluster data.111In principle, our method can uncover skills that are not describable through these templates, but templating allows automatic labeling.
+Our templating approach is similar to the automated explainer tool N2G (Foote et al., 2023), a popular ngram-based evaluation metric for unsupervised interpretability methods (Gao et al., 2024).
+Our automatic labels are based on the target token and its preceding trigram.
+To obtain automatic labels for a cluster, we compute the frequency of each POS tag (tagged by spacy) in it.
+We then automatically label each cluster with the smallest set of POS tags required to compose of the cluster’s 4-gram samples.
+For example, if over of the token instances in a cluster are preceded by a comma <,>, the cluster would be automatically labeled as 1 token after PUNCT.
+We consider basis vectors with at least one cluster with a simple label—one including at most two POS tags (not counting <PAD> tokens).
+Starting with 30 basis vectors, we remove 4 because the average decomposed loss increases.
+We find of the remaining basis vectors have at least one cluster with a simple label as defined.
+We refine these automatically assigned labels by examining the most frequent tokens in the cluster and the ten examples closest to the cluster medoid and manually selecting a label that is consistent with the automatic POS label and the ten closest examples to the medoid.
+See Appendix E.2 for further labeling details.
 5.2 Results
 On language models, POLCA clustering again reveals hidden breakthroughs along each basis vector.
-To analyze the learned skills represented by each cluster, we look for syntactic and lexical patterns shared by the cluster data.
-After tagging tokens by part-of-speech (POS) using spacy, we label each cluster by computing its most common token and POS tag sequences.
-We confirm these automatically assigned cluster labels by examining the top 10 contexts closest to the medoid of the cluster.
-We report a selection of clusters from specific basis vectors in Figure 6 (see Appendix H for more examples).
-Using POLCA clustering on projected loss trajectories for our basis vectors, we find token subpopulations that correspond to various grammatical constructions and have breakthroughs in the loss projected onto that basis.
-For instance, we show a cluster corresponding to predicting < to> and < from> after the first clause in a sentence in Figure 6a(6ii).
+We show a selection of clusters from specific basis vectors in Figure 6 and others in Appendix K (Figures 12 and 14).
+Clustering projected loss trajectories along each basis vector, we find groups that correspond to various grammatical constructions.
+Some show apparent breakthroughs in their projected loss, like the cluster corresponding to predicting < to> and < from> after the first clause in a sentence (Figure 6a(6i)).
 We also observe clusters whose projected loss trajectories move in opposing directions along certain basis vectors.
-For instance, in Figure 6b(6i) and 6b(6ii), cluster 1 contains appositive noun phrases and cluster 2 has noun phrases with similar syntactic patterns that are not appositive (such as list items).
-These clusters visibly mirror each other’s movement in direction, if not in magnitude—decreases in loss are generally larger than the opposing cluster’s increases in loss which mirror them.
+For instance, in Figure 6b, cluster 1 contains appositive noun phrases and cluster 2 has syntactically similar—but non-appositive—noun phrases (such as list items).
+These clusters visibly mirror each other’s movement—though the clustered decreases in loss are generally larger than their opposing cluster’s increases in loss.
 Figure 6 shows that despite their smooth exact loss curves, POLCA clusters have sudden changes in their decomposed loss curves at different points during training.
@@
 We conclude that POLCA reveals breakthroughs in the decomposed loss that are obscured in the exact loss.
-Through clustering, POLCA can explain how different skills are learned during language model training.
+Through clustering, POLCA can explain how different skills are learned during training.
 6 Conclusions
@@
 These are promising findings for meaningfully interpreting large models.
-By recovering breakthroughs in identifiable skills, we support the hypothesis that high-dimensional learning typically entails a series of phase transitions at various scales.
-When a phase transition appears in training, it suggests a naturally discrete category; the model either knows the concept or doesn’t know it, with little middle ground.
-Humans think in categorical concepts, so they are far more interpretable than the continuous data interpolations that appear in much of learning theory.
-Limitations
-Our method of constructing a basis is inspired by the existing literature on training in restricted subspaces, but represents an obvious site of improvement.
-The top eigenvectors of the Hessian, like the axis-aligned basis, could represent many concepts in superposition.
-Therefore, some non-orthogonal basis might represent interpretable concepts more cleanly than our orthogonal basis, though it would no longer provide a low-rank decomposition.
-Furthermore, our basis is constructed by local curvature and then filtered to favor directions of long-term movement; other bases may favor long-term movement by construction.
-In general, we consider the ideal basis to be an open question.
-We use small models so that we can have very fine-grained checkpoints for POLCA computation.
-Scaling this work to larger models may require using a basis that is less computationally expensive to compute than Hessian eigenvectors, but our results from Appendix E indicate that this is likely possible with limited impact on the cluster quality.
+By recovering breakthroughs in identifiable skills, we support the hypothesis that high-dimensional learning typically entails a series of breakthroughs at various scales.
+When a breakthrough appears in training, it suggests a naturally discrete category; the model either knows the concept or doesn’t know it, with little middle ground.
+Humans think in categorical concepts, so these breakthroughs are far more interpretable than the continuous data interpolations often assumed in learning theory.
+Reproducibility statement
+We implement our models and experiments using open-source libraries and datasets.
+We provide detailed hyperparameters in Appendix D and a thorough explanation of the experimental setup in Section E.
+Our code is available at https://github.com/skangasl/POLCA.
+Ethics statement
+This work provides a method for better understanding the training dynamics of language models. The trained models may contain biases from the training datasets.
 Acknowledgements
@@
 This work was enabled in part by a gift from the Chan Zuckerberg Initiative Foundation to establish the Kempner Institute for the Study of Natural and Artificial Intelligence.
-Appendix A Derivation of approximate second-order term
+Appendix A Limitations and future work
+Our method of constructing a basis is inspired by the existing literature on training in restricted subspaces, but represents an obvious site of improvement.
+The top eigenvectors of the Hessian, like the axis-aligned basis, could represent many concepts in superposition.
+Therefore, some non-orthogonal basis might represent interpretable concepts more cleanly than our orthogonal basis, though it would no longer provide a low-rank decomposition.
+Furthermore, our basis is constructed by local curvature and then filtered to favor directions of long-term movement; other bases may favor long-term movement by construction.
+In general, we consider the ideal basis to be an open question.
+Our experiments are limited to small models.
+The two main challenges with scaling POLCA are the Hessian basis computation and the frequency of checkpoints used to sample the POLCA trajectories.
+Scaling this work to larger models may require using a basis that is less computationally expensive to compute than Hessian eigenvectors, but our results from Appendix H indicate that this is likely possible with limited impact on the cluster quality.
+The small scale of models that we use in our experiments allows for very high granularity of checkpoints used to compute both the basis and the POLCA trajectories.
+For larger models, this may be computationally infeasible and a lower checkpoint frequency may be needed, resulting in less signal for clustering the POLCA trajectories.
+Our current experiments are limited to language models.
+However, in principle our approach is model-agnostic and can be applied to any deep neural network.
+Applying POLCA to other modalities is an exciting direction for future work.
+The labeling approach that we use in the natural language setting relies on POS tagging.
+This labeling strategy allows for unsupervised, automatic identification of these syntactic skills and ensures strict interpretable labels.
+However, it fails to capture many human-interpretable language modeling skills.
+The discarded vectors may (and likely do) contain other interpretable skill clusters that are not found by automated labeling.
+Appendix B Statement on use of large language models
+We used generative AI for debugging and minor grammar edits when writing. The authors made all significant contributions to the research, analyses, and writing.
+Appendix C Derivation of approximate second-order term
 We can approximate the difference between the gradient at time and as
@@
 If this assumption is violated, then the scaling factor in the following second-order Taylor term will be minuscule on the datapoint level.
-In practice, we have found that the second-order term has limited impact at the datapoint level (see Appendix F), but we nonetheless use it to improve our approximation.
+In practice, we have found that the second-order term has limited impact at the datapoint level (see Appendix I), but we nonetheless use it to improve our approximation.
 Then we may approximate it as:
-Appendix B Hyperparameters
+Empirical Validation
+We run a small-scale empirical study to test the accuracy of this estimation.
+For a sample of tokens and the first POLCA checkpoints and the top basis vectors in the arithmetic setting, we compute the root mean squared error (RMSE) between the approximated second-order term and the true second-order term averaged across tokens, checkpoints, and basis vectors.
+We find that our approximation has an RMSE of when compared to the ground truth second-order term, indicating that this approximation is close to the real value.
+Appendix D Hyperparameters
 In the tables below, we provide the hyperparameters used during training of the models in the synthetic arithmetic and language modeling settings.
 We use NVIDIA H100 80GB HBM3 GPUs for our experiments and run each training run on a single GPU.
+We selected the clustering hyperparameters to maximize empirical performance in the synthetic setting and used similar values relative to the number of tokens in the natural language experiments.
+We chose the POLCA hyperparameters to maximize the frequency of POLCA checkpoints and basis checkpoints within computational constraints.
 | Hyperparameter | Value |
 | Number of Parameters | 9475594 |
-| Iterations | 10000 |
+| Steps | 10000 |
 | Epochs | 1 |
@@
 | Hyperparameter | Value |
+| Basis checkpoint interval (steps) | 200 |
+| T | 50 |
+| k | 1 |
+| POLCA checkpoint interval (steps) | 5 |
+| Hyperparameter | Value |
+| Clustering algorithm | HDBSCAN |
+| Minimum cluster size | 150 |
+| Minimum samples | Number of tokens / 15 |
+| Hyperparameter | Value |
 | Number of Parameters | 40274737 |
-| Iterations | 14000 |
+| Steps | 14000 |
 | Epochs | 1 |
@@
 | LR Schedule |
-Appendix C Exact loss trajectories for the digit and carry skills
-Appendix D Decomposition strategy comparison
-We investigate whether POLCA is required for decomposing the loss. We test three strategies:
-Loss: Exact loss
-Fisher information: We approximate the empirical Fisher Information as as in [2]. For each basis vector , the Fisher Information projected onto is .
+| Hyperparameter | Value |
+| Basis checkpoint interval (steps) | 750 |
+| T | 30 |
+| k | 1 |
+| POLCA checkpoint interval (steps) | 200 |
+| Hyperparameter | Value |
+| Clustering algorithm | HDBSCAN |
+| Minimum cluster size | 300 |
+| Minimum samples | Number of tokens / 15 |
+Appendix E Experimental details
+E.1 Arithmetic setting
+Setup.
+In the arithmetic experiments in Section 4, we train a 3-layer (9m parameter) Transformer model with embedding dimension 512, 4 attention heads per layer, and an MLP dimension of 2048 (Nanda & Bloom, 2022).
+For a validation set with 1250 data points and 5000 output tokens, we compute the loss and POLCA values for each token at intervals of 20 steps throughout training.
+We compute the POLCA basis using the eigenvectors of the Hessian estimated using a 1250 data point sample of the training set.
+We compute a new basis vector every 200 steps for a total of 50 basis vectors.
+Labeling details.
+We automatically label each token with the ground truth value of the digit and carry skills using the definition of these two skills.
+We label the digit skill based on the position of the token in the output: the first token is 1000s, the second token is 100s, the third is 10s, and the fourth is 1s.
+We label the carry skill by computing the sum up to the next lowest digit place and determining whether it resulted in a carry to the current token.
+For instance, for an output in the 10s place, if the two inputs in the 1s place sum to 10 or higher, then it will be labeled with "carry", otherwise it will be labeled with "no carry".
+E.2 Natural language setting
+Setup.
+For the natural language experiments in Section 5, we train on the English Wikipedia dataset (Wikimedia Foundation, 2022) from March 2022.
+We use the same POLCA setup as in the arithmetic addition setting.
+We train a 3-layer (40m parameter) transformer model with embedding dimension 512, 4 attention heads, and an MLP dimension of 2048 (Nanda & Bloom, 2022).
+We compute the loss and POLCA values for each token on a validation set of 12600 output tokens.
+We analyze intermediate checkpoints at intervals of 200 steps throughout training.
+We apply POLCA to the basis derived from the eigenvectors of the Hessian estimated using a 1000 data point sample of the training set as detailed in Algorithm 1 with .
+We compute a new basis vector every 750 steps.
+Labeling details.
+We label each token and the three tokens before it using spacy for part-of-speech (POS) tagging.
+This produces a sequence of four POS tags as the automatic POS label for each token.
+To label a given cluster, we compute the frequency (across the tokens in the cluster) of each POS tag at each index.
+For each index in the sequence, we then automatically label the cluster with the smallest set of POS tags at that index required to make up 70% of the cluster.
+We then filter out any labels that require more than 2 POS tags to describe the cluster.
+To generate the labels reported in Section 5, we manually refine the automatically generated label by looking at the top 10 contexts closest to the medoid of the cluster.
+Although these manually refined labels are challenging to verify automatically, we ensure that the contexts closest to the medoid follow the manual label and that the manual label follows the automatic label generated using the POS tags.
+Appendix F Exact loss trajectories for the digit and carry skills
+Appendix G Decomposition strategy comparison
+We investigate whether POLCA is required for decomposing the loss.
+We expand on the results from Table 1 by showing the top three homogeneity scores for the carrying skill and adding empirical Fisher information and first order POLCA results.
+To compute the homogeneity score for a given basis vector, we take the maximum fraction of carries in any given cluster for that basis vector (or across all of the clusters for exact loss or change in exact loss).
+We then report the top three homogeneities across the full basis, as well as the recall and F1 score for the corresponding clusters.
+We note that we exclude HDBSCAN outliers from the recall and F1 computations.
+Importantly, we only consider clusters for which over of the tokens with the carry skill correspond to the s or s place, since carrying to the s place corresponds to simply predicting a or in the first position (see Figure 3 for reference) and is recovered with high homogeneity for all trajectory types except exact loss.
+We compare carry skill homogeneity across the following trajectory types:
+Loss: Exact loss trajectories.
+Change in exact loss: We compute the change in exact loss by subtracting the loss at checkpoint from the loss at checkpoint for each timestep in the exact loss trajectory.
+Fisher information: We approximate the empirical Fisher Information as as in Achille et al. (2017). For each basis vector , the Fisher Information projected onto is .
+Loss Change Allocation (LCA): We compute the datapoint-wise LCA trajectories (Equation 1) projected onto the parameters that have the top highest magnitudes at the end of training.
+First order POLCA: We calculate the POLCA trajectory without the second order term (Equation 2)
 POLCA: We compute the projected loss (Equation 7).
-While using the projected gradient is less computationally intensive than computing the entire POLCA term, the results from Table 3 demonstrate that clustering on solely the Fisher information does not yield sufficiently homogeneous carry clusters.
-| Decomposition strategy | Top vector skill homogeneities | ||
-| 1 | 2 | 3 | |
-| Loss | 0.514 | - | |
-| Fisher information | 0.664 | 0.643 | 0.637 |
-| POLCA | 0.973 | 0.946 | 0.903 |
-Appendix E POLCA basis comparison
+The results from Table 8 demonstrate that POLCA finds the most homogeneous clusters with respect to the carrying skill.
+Moreover, POLCA and first order POLCA have comparable F1 scores.
+Note that many low homogeneity clusters can have high recall because they are large and thus contain most of the carry tokens while having low homogeneity.
+| Decomposition strategy | Cluster | |||
+| Number | Homogeneity | Recall | F1 | |
+| Loss | 1 | 0.514 | 0.771 | 0.617 |
+| Change in exact loss | 1 | 0.524 | 0.958 | 0.678 |
+| 1 | 0.664 | 0.947 | 0.781 | |
+| Fisher information | 2 | 0.643 | 0.740 | 0.688 |
+| 3 | 0.637 | 0.874 | 0.737 | |
+| 1 | 0.792 | 0.772 | 0.782 | |
+| Loss change allocation (LCA) (Lan et al., 2020) | 2 | 0.614 | 0.873 | 0.721 |
+| 3 | 0.592 | 0.626 | 0.609 | |
+| 1 | 0.948 | 0.767 | 0.848 | |
+| First order POLCA | 2 | 0.928 | 0.769 | 0.841 |
+| 3 | 0.887 | 0.751 | 0.813 | |
+| 1 | 0.973 | 0.736 | 0.838 | |
+| POLCA | 2 | 0.946 | 0.773 | 0.850 |
+| 3 | 0.903 | 0.762 | 0.827 | |
+We also compare the fraction of clusters with hidden breakthroughs for each type of trajectory.
+To compute whether or not a given cluster has a hidden breakthrough, we use Equation 10 with to identify breakthroughs past step where the exact loss plateaus.
+We find that POLCA produces the highest fraction of clusters with hidden breakthroughs.
+We hypothesize that this is mostly due to the basis construction, since the Fisher information and first order POLCA (both of which are computed using the same basis as POLCA) produce the next highest fraction of clusters with hidden breakthroughs.
+| Decomposition strategy | Fraction of clusters with hidden breakthroughs |
+| Loss | 0.0 |
+| Change in exact loss | 0.0 |
+| Fisher information | 0.284 |
+| Loss change allocation (LCA) (Lan et al., 2020) | 0.019 |
+| First order POLCA | 0.307 |
+| POLCA | 0.355 |
+Appendix H POLCA basis comparison
 We test ablated bases to analyze the effect of basis choice on the POLCA breakthrough clustering.
@@
 Top Hessian eigenvectors: basis computed using Algorithm 1
-We find in Table 4 that these ablations result in only slightly less homogeneous carry clusters, indicating that different bases can be used for larger experiments to trade off between compute and interpretability.
-| POLCA basis | Top vector skill homogeneities | ||
-| 1 | 2 | 3 | |
-| Random orthonormal | 0.902 | 0.858 | 0.838 |
-| Random shuffled Hessian | 0.856 | 0.852 | 0.834 |
-| Top Hessian eigenvectors | 0.973 | 0.946 | 0.903 |
-Appendix F Second versus first order POLCA approximation
+We find in Table 10 that these ablations result in only slightly lower quality clusters with respect to homogeneity (although random orthonormal vectors have lower recall than the other two bases on average), indicating that different bases can be used for larger experiments to trade off between compute and interpretability.
+We also compute the fraction of clusters with hidden breakthroughs for each basis in Table 11 and find that the random orthonormal basis has a significantly lower fraction of hidden breakthroughs recovered than the other two approaches, indicating that this random orthonormal basis is not sufficient to find hidden breakthroughs late in training.
+In addition to these bases, we have tested a variety of additional basis constructions, such as a stacked Jacobian, Hessian computed using a sliding window, and Hessian computed at the end of training, and chose to use the top Hessian eigenvectors since they had the best performance in the arithmetic setting.
+| POLCA basis | Cluster | |||
+| Number | Homogeneity | Recall | F1 | |
+| 1 | 0.902 | 0.655 | 0.759 | |
+| Random orthonormal | 2 | 0.858 | 0.533 | 0.658 |
+| 3 | 0.838 | 0.586 | 0.689 | |
+| 1 | 0.856 | 0.699 | 0.769 | |
+| Random shuffled Hessian | 2 | 0.852 | 0.734 | 0.789 |
+| 3 | 0.834 | 0.730 | 0.779 | |
+| 1 | 0.973 | 0.736 | 0.838 | |
+| POLCA | 2 | 0.946 | 0.773 | 0.850 |
+| 3 | 0.903 | 0.762 | 0.827 | |
+| POLCA basis | Fraction of clusters with hidden breakthroughs |
+| Random orthonormal | 0.031 |
+| Random shuffled Hessian | 0.304 |
+| POLCA | 0.355 |
+Appendix I Second versus first order POLCA approximation
 | Cosine similarity | L2 norm |
 | 5.4891 E-4 | 0.99987 |
-Appendix G Additional arithmetic language modeling clusters
-Appendix H Additional natural language clusters
+Appendix J Additional arithmetic language modeling clusters
+Appendix K Additional natural language clusters
 | Cluster | Label | Contexts closest to medoid |
@@
 […]
+| Cluster | Label | Contexts closest to medoid |
+| 1 | Punctuation after noun phrase | Pao Newspaper Company Limited () under Ho Man-fat.
+It was initially published every three days, in the 6th and 5th centuries BC, to denote some type of a sibant sound, president of the United States the commander-in-chief of the United States Armed Forces.
+Many presidents, |
+[…]
+| 3 |
+< of>
+official from the Electorate of the Palatinate who served Brandenburg-Prussia.
+He was the son of and civil parish in Milton Keynes, ceremonial Buckinghamshire, England.
+For local government purposes, it is part of ering as it extends outward from the coast.
+Cape Lookout State Park is located on the north side of |
 | Cluster | Label | Contexts closest to medoid |
@@
 […]
+| Cluster | Label | Contexts closest to medoid |
+| 1 |
+Token after <,>
+[…]
+| 2 | Repeated <PAD> tokens | <PAD><PAD><PAD><PAD><PAD><PAD><PAD><PAD><PAD> <PAD><PAD><PAD><PAD><PAD><PAD><PAD><PAD><PAD> <PAD><PAD><PAD> <PAD><PAD><PAD><PAD><PAD><PAD><PAD><PAD><PAD> <PAD><PAD><PAD><PAD><PAD><PAD><PAD><PAD><PAD> <PAD><PAD><PAD> <PAD><PAD><PAD><PAD><PAD><PAD><PAD><PAD><PAD> <PAD><PAD><PAD><PAD><PAD><PAD><PAD><PAD><PAD> <PAD><PAD><PAD> |